# Advancing GPU IPC for Stiff Affine-Deformable Simulation

### v2

@@ Abstract.
Incremental Potential Contact (IPC) is a widely used, robust, and accurate method for simulating complex frictional contact behaviors. However, achieving high efficiency remains a major challenge, particularly as material stiffness increases, which leads to slower Preconditioned Conjugate Gradient (PCG) convergence, even with the state-of-the-art preconditioners. In this paper, we propose a fully GPU-optimized IPC simulation framework capable of handling materials across a wide range of stiffnesses, delivering consistent high performance and scalability with up to 10$\times$ speedup over state-of-the-art GPU IPC methods. Our framework introduces three key innovations: 1) A novel connectivity-enhanced Multilevel Additive Schwarz (MAS) preconditioner on the GPU, designed to efficiently capture both stiff and soft elastodynamics and improve PCG convergence at a reduced preconditioning cost. 2) A $C^{2}$-continuous cubic energy with an analytic eigensystem for strain limiting, enabling more parallel-friendly simulations of stiff membranes, such as cloth, without membrane locking. 3) For extremely stiff behaviors where elastic waves are barely visible, we employ affine body dynamics (ABD) with a hash-based multi-layer reduction strategy for fast Hessian assembly and efficient affine-deformable coupling. We conduct extensive performance analyses and benchmark studies to compare our framework against state-of-the-art methods and alternative design choices. Our system consistently delivers the fastest performance across soft, stiff, and hybrid simulation scenarios, even in cases with high resolution, large deformations, and high-speed impacts. Our framework will be fully open-sourced upon acceptance.
@@ 1. Introduction
In summary, there are still plenty of rooms for optimizing linear solver preconditioners, strain limiting, global Hessian matrix assembly, etc., for realizing a highly GPU-optimized IPC framework that can efficiently simulate large-scale affine-deformable coupled scenarios. In this paper, we propose such a framework, achieving up to 10×10$\times$10 × speedup compared to GIPC via the following 3 major innovations:
@@
  • A $C^{2}$-continuous cubic strain-limiting energy with an analytic eigensystem, enabling realistic cloth simulation without membrane locking (section 5). As numerical eigendecomposition and line search filtering for the feasibility of the strain limits are not needed, our model supports highly GPU-parallelized computations.
  • A hash-based multi-layer reduction strategy for fast Hessian matrix assembly (section 6). Our strategy significantly reduces the number of numerical operations, and it enables the development of a memory-efficient symmetric blockwise sparse matrix-vector multiplication method to further boost PCG performance.
@@ 2. Related Works
One of the most critical design choices on the GPU is to apply iterative solvers instead of direct solvers for linear systems. Multigrid is a popular GPU-friendly preconditioner often used in iterative linear solvers, such as the PCG method, for efficiently solving ill-conditioned systems. The major challenge in multigrid methods is the development of effective restriction and prolongation operators, especially for unstructured meshes. Researchers in computer graphics have explored various geometric multigrid methods (GMG) (Wang et al., 2020, 2018; Xian et al., 2019; Zhu et al., 2010) by constructing mesh hierarchies, where numerous implementation issues are often encountered. Compared to GMG, another type of multigird, the algebraic multigrid (AMG) methods (Naumov et al., 2015; Demidov, 2019; Tamstorf et al., 2015), do not require explicitly constructing coarser geometries. However, they often require a significant amount of precomputation to analyze the structure of the matrix, making them more expensive when handling problems with varying matrix structures. This inspires Wu et al. (2022) to propose a multilevel additive schwarz (MAS) (Dryja and Widlund, 1990) preconditioner, incorporating the idea of domain decomposition to achieve fast convergence with low per-iteration costs when applied to PCG. But its node reordering based on Morton code is not aware of the mesh connectivity, which may lead to suboptimal aggregation. We thus propose a connectivity-enhanced MAS based on METIS (Karypis and Kumar, 1998) to tackle this issue and achieve even faster performance.
@@
To realistically simulate cloth using linear elements, strain limiting is often applied to avoid membrane locking. Most existing methods (Narain et al., 2012; Wang et al., 2010; Thomaszewski et al., 2009; English and Bridson, 2008; Goldenthal et al., 2007) enforce a bound constraint on the strain measure per element using augmented Lagrangian approaches or post-projection techniques. While these methods do not guarantee strict satisfaction of strain limits, they can effectively simulate membrane locking-free behaviors. However, when coupled with frictional contact, these methods often introduce artifacts due to the need for smaller time steps or difficulties in achieving solver convergence. More recently, Li et al. (2021) applied a barrier method to handle strain-limiting constraints, guaranteeing exact constraint satisfaction in a monolithic manner. However, it requires expensive backtracking-based line search filtering and numerical eigendecomposition, which are both GPU-unfriendly. Here, the backtracking line search is responsible for finding configurations that are both feasible and with a smaller objective function. With the development of analytical eigenanalysis on distortion energies, such as the isotropic (Smith et al., 2019; Lin et al., 2022) and anisotropic (Kim et al., 2019) elasticity, membrane (Panetta, 2020; Kim, 2020), bending (Wang et al., 2023b; Wu and Kim, 2023), and contact (Shi and Kim, 2023; Huang et al., 2024) energies, they have become popular components of modern GPU simulation methods. We thus propose a cubic strain-limiting energy with analytic eigensystems so that realistic cloth dynamics can be more efficiently simulated without the need of numerical eigendecomposition and backtracking-based line search filtering.
@@ 3. Background and preliminaries
Figure 2. An overview of 1 Newton iteration in our unified GPU IPC framework. Our simulation focuses on cloth, soft bodies with varying stiffness, and stiff affine bodies (Lan et al., 2022a). Cloth and soft bodies are modeled using FEM. Contact between all objects is handled using IPC (Li et al., 2020a). We solve the system with a Newton-PCG solver and use Continuous Collision Detection (CCD) to filter the search direction and prevent penetration, applying backtracking line search to ensure energy decrease. Our contributions, highlighted in light pink blocks, include a novel strain limiting energy with an analytic eigensystem for efficient PSD projection, an optimized linear solver with a connectivity-enhanced MAS preconditioner, and a highly-parallelized global Hessian assembly strategy for affine-deformable coupling.
@@ 3.1. Unified Incremental Potential
where $\mathbf{q}\in\rm I\!R^{12\alpha}$ is the reduced space coordinates of the $\alpha$ affine bodies. For affine body $j$, $\mathbf{q}_{j}=[\mathbf{p}_{j}^{T},\mathbf{A}_{j1}^{T},\mathbf{A}_{j2}^{T},\mathbf{A}_{j3}^{T}]^{T}$, with $\mathbf{p}_{j}\in\rm I\!R^{3}$ the translation vector and $\mathbf{A}_{j}\in\rm I\!R^{3\times 3}$ the affine deformation matrix. Here, $\mathbf{x}_{i}(\mathbf{q}_{j})=\mathbf{A}_{j}\bar{\mathbf{x}}_{i}+\mathbf{p}_{j}$ is the current position of full-space node $i$ on affine body $j$ for measuring contact energies, with $\bar{\mathbf{x}}_{i}$ the rest-state full space coordinates. The Jacobian matrix $\mathbf{J}_{ij}=\frac{\partial\mathbf{x}_{i}}{\partial\mathbf{q}_{j}}\in\rm I\!R^{3\times 12}$ maps information between the reduced and full spaces. For example, the mass matrix $\mathbf{M}_{r}$ of affine body $j$ is calculated as $\mathbf{M}_{r}=\sum_{i}m_{i}\mathbf{J}_{ij}^{T}\mathbf{J}_{ij},$ where $m_{i}$ is mass of each node $i$. Similarly, the contact gradient of affine body $j$ is calculated as $\nabla_{\mathbf{q}_{j}}B=\sum_{i}\mathbf{J}_{ij}^{T}\nabla_{\mathbf{x}_{i}}B.$ From this, we can see that computing contact forces and Hessian matrices for the reduced DOFs could require accumulating values from a significant amount of contact pairs, which can be expensive.
@@
To accelerate local Hessian computation on the GPU, we utilize potential energies with analytic eigensystems (subsection 3.2). Specifically, for collision energy $B$ and friction energy $D$, we employ the eigenanalysis from GIPC (Huang et al., 2024). The Stable Neo-Hookean model (Smith et al., 2018) is applied for $\Psi_{vol}$, while the bending energy from (Wu and Kim, 2023) is used for $\Psi_{bend}$. To efficiently assemble $\Psi_{memb}$ and $\Psi_{strain}$, we use Kim (2020)’s membrane energy and propose a cubic strain-limiting energy with analytic eigensystem, detailed in section 5.
@@ 3.2. IP Minimization and Linear Solves
To minimize $E(\mathbf{u})$, where $\mathbf{u}=\{\mathbf{q},\mathbf{x}\}$, line search methods are often used, which iteratively search along a descent direction by minimizing a local quadratic proxy of the IP in iteration $i$:
(6) $E_{i}(\mathbf{u})=E(\mathbf{u}_{i})+\left(\mathbf{u}-\mathbf{u}_{i}\right)^{T}\nabla E(\mathbf{u}_{i})+\frac{1}{2}\left(\mathbf{u}-\mathbf{u}_{i}\right)^{T}\mathbf{P}(\mathbf{u}_{i})\left(\mathbf{u}-\mathbf{u}_{i}\right).$
The descent direction is obtained by solving the linear system $\mathbf{P}(\mathbf{u}_{i})\mathbf{d}=-\nabla E(\mathbf{u}_{i})$, where $\mathbf{d}=\mathbf{u}_{i}^{*}-\mathbf{u}_{i}$ and $\mathbf{u}_{i}^{*}=\operatorname*{arg\,min}_{\mathbf{u}}E_{i}(\mathbf{u})$. Here, $\mathbf{P}$ is a symmetric positive definite (SPD) proxy matrix approximating $\nabla^{2}E$ for fast convergence, with projected Newton (Li et al., 2020a) or Gauss-Newton (Huang et al., 2024) approximations shown effective. The new iterate is computed as $\mathbf{u}_{i+1}=\mathbf{u}_{i}+\alpha\mathbf{d}$, where $\alpha$ is the step size calculated via backtracking line search, with filtering to ensure feasibility imposed by contact and strain limits (Li et al., 2021). The minimization terminates when the iterate is sufficiently close to the local minimum, e.g. when $\|\mathbf{d}\|\leq\varepsilon_{d}$, where $\varepsilon_{d}$ is the Newton tolerance.
@@ 4.1. Discussions of MAS
MAS constructs the domain hierarchy by first sorting all mesh vertices using Morton codes, and then grouping these vertices into subdomains of size $N$. Next, connected vertices within each subdomain are merged to create the next-level simulation domain. These merged vertices, called super nodes, are then grouped again to form the domains at subsequent levels until a maximum level is reached, no further merges could be performed, or only a single subdomain is present. During preconditioning, the inverse of each subdomain’s Hessian matrix is multiplied with the mapped input vector, and the resulting vectors are mapped and summed to produce the output. With a fixed number of levels, maximizing node connectivity within each subdomain is critical to making dimensionality reduction more effective, which allows information to propagate farther at coarser levels and ensures faster convergence. See our supplemental document for more details.
@@ 4.2. Connectivity-Enhanced MAS Construction
To implement this approach, we first construct a mapping array of size $N\cdot M$, which maps each node to a slot in its corresponding MAS subdomain, as shown in Figure 7. This mapping array is then used to build the MAS preconditioner. To ensure correct data access during construction, we also create a remapping array, which helps GPU threads access the correct node data for each entry in the mapping array. Padded entries, where no nodes are mapped, are marked with the value $-1$ (see Figure 7). The remaining MAS levels are constructed using the original method based on node connectivity. By updating the mapping per time step considering the connectivity introduced by the current contact pairs, our approach reduces connectivity issues and enables more effective aggregation, improving the convergence speed of the MAS-preconditioned linear solver by up to 2.4$\times$.

### v3

# StiffGIPC: Advancing GPU IPC for Stiff Affine-Deformable Simulation

@@ Abstract.
Incremental Potential Contact (IPC) is a widely used, robust, and accurate method for simulating complex frictional contact behaviors. However, achieving high efficiency remains a major challenge, particularly as material stiffness increases, which leads to slower Preconditioned Conjugate Gradient (PCG) convergence, even with the state-of-the-art preconditioners. In this paper, we propose a fully GPU-optimized IPC simulation framework capable of handling materials across a wide range of stiffnesses, delivering consistent high performance and scalability with up to 10$\times$ speedup over state-of-the-art GPU IPC methods. Our framework introduces three key innovations: 1) A novel connectivity-enhanced Multilevel Additive Schwarz (MAS) preconditioner on the GPU, designed to efficiently capture both stiff and soft elastodynamics and improve PCG convergence at a reduced preconditioning cost. 2) A $C^{2}$-continuous cubic energy with an analytic eigensystem for inexact strain limiting, enabling more parallel-friendly simulations of stiff membranes, such as cloth, without membrane locking. 3) For extremely stiff behaviors where elastic waves are barely visible, we employ affine body dynamics (ABD) with a hash-based two-level reduction strategy for fast Hessian assembly and efficient affine-deformable coupling. We conduct extensive performance analyses and benchmark studies to compare our framework against state-of-the-art methods and alternative design choices. Our system consistently delivers the fastest performance across soft, stiff, and hybrid simulation scenarios, even in cases with high resolution, large deformations, and high-speed impacts. Our framework will be fully open-sourced upon acceptance.
@@ 1. Introduction
In summary, there is still plenty of room for optimizing linear solver preconditioners, strain limiting, global Hessian matrix assembly, etc., for realizing a highly GPU-optimized IPC framework that can efficiently simulate large-scale affine-deformable coupled scenarios. In this paper, we propose such a framework, achieving up to 10×10$\times$10 × speedup compared to GIPC via the following 4 major innovations:
@@
  • A $C^{2}$-continuous cubic inexact strain-limiting energy with an analytic eigensystem, enabling realistic cloth simulation without membrane locking (section 5). As numerical eigendecomposition and line search filtering for the feasibility of the strain limits are not needed, our model supports highly GPU-parallelized computations.
  • A hash-based two-level reduction strategy for fast Hessian matrix assembly in affined-deformable coupled systems (section 6). Our strategy significantly reduces the number of numerical operations, and it enables the development of a memory-efficient symmetric blockwise Sparse Matrix-Vector Multiplication (SpMV) method to further boost PCG performance.
@@ 2. Related Works
One of the most critical design choices on the GPU is to apply iterative solvers instead of direct solvers for linear systems. Multigrid is a popular GPU-friendly preconditioner often used in iterative linear solvers, such as the PCG method, for efficiently solving ill-conditioned systems. The major challenge in multigrid methods is the development of effective restriction and prolongation operators, especially for unstructured meshes. Researchers in computer graphics have explored various geometric multigrid methods (GMG) (Wang et al., 2020, 2018; Xian et al., 2019; Zhu et al., 2010) by constructing mesh hierarchies, where numerous implementation issues are often encountered. Compared to GMG, another type of multigrid, the algebraic multigrid (AMG) methods (Naumov et al., 2015; Demidov, 2019; Tamstorf et al., 2015), do not require explicitly constructing coarser geometries. However, they often require a significant amount of precomputation to analyze the structure of the matrix, making them more expensive when handling problems with varying matrix structures. This inspires Wu et al. (2022) to propose a multilevel additive schwarz (MAS) (Dryja and Widlund, 1990) preconditioner, incorporating the idea of domain decomposition to achieve fast convergence with low per-iteration costs when applied to PCG. But its node reordering based on Morton code is not aware of the mesh connectivity, which may lead to suboptimal aggregation. We thus propose a connectivity-enhanced MAS based on METIS (Karypis and Kumar, 1998) to tackle this issue and achieve even faster performance.
Sparse Matrix-Vector Multiplication
Optimizing sparse matrix-vector multiplication (SpMV), the primary bottleneck in iterative linear solvers, is as crucial as accelerating the convergence. Gkountouvas et al. (2013) introduced a symmetric variant of the Compressed Sparse eXtended (CSX) (Kourtis et al., 2011) format, leveraging nonzero indexing compression to reduce memory traffic in symmetric SpMV. Eberhardt and Hoemmen (2016) proposed optimizations using the Block Compressed Row Storage (BCSR) format, while GIPC (Huang et al., 2024) developed a matrix-free SpMV method tailored to their PCG solver. A recent survey on Sparse Matrix-Matrix Multiplication (SpMM) (Gao et al., 2023) further discusses relevant optimization strategies. Currently, cuSPARSE [^1] https://developer.nvidia.com/cusparse provides the fastest SpMV implementation via the Block Sparse Row (BSR) format, utilizing GPU-optimized instructions and warp reduction. However, taking full advantage of BSR SpMV requires an efficient Hessian assembly strategy, such as Tang et al. (2016). Furthermore, since SpMV is memory-bound and the Hessian is symmetric, leveraging symmetry can enhance performance, but cuSPARSE does not support this optimization. To address these limitations, we introduce a fast Hessian assembly method using a sorted block coordinate format, enabling efficient warp reduction and symmetry exploitation, surpassing cuSPARSE BSR SpMV in performance.
@@
To realistically simulate cloth using linear elements, strain limiting is often applied to avoid membrane locking. Most existing methods (Narain et al., 2012; Wang et al., 2010; Thomaszewski et al., 2009; English and Bridson, 2008; Goldenthal et al., 2007) enforce a bound constraint on the strain measure per element using augmented Lagrangian approaches or post-projection techniques. While these methods do not guarantee strict satisfaction of strain limits, they can effectively simulate membrane locking-free behaviors. However, when coupled with frictional contact, these methods often introduce artifacts due to the need for smaller time steps or difficulties in achieving solver convergence. More recently, Li et al. (2021) applied a barrier method to handle strain-limiting constraints, guaranteeing exact constraint satisfaction in a monolithic manner. However, it requires expensive backtracking-based line search filtering and numerical eigendecomposition, which are both GPU-unfriendly. Here, two backtracking line searches are needed in each Newton iteration: the first, often called line search filtering, determines the maximum step size that maintains a feasible configuration, while the second starts from this step size and further refines it to find a configuration with a lower objective function. With the development of analytical eigenanalysis on distortion energies, such as the isotropic (Smith et al., 2019; Lin et al., 2022) and anisotropic (Kim et al., 2019) elasticity, membrane (Panetta, 2020; Kim, 2020), bending (Wang et al., 2023b; Wu and Kim, 2023), and contact (Shi and Kim, 2023; Huang et al., 2024) energies, they have become popular components of modern GPU simulation methods. We thus propose a cubic inexact strain-limiting energy with analytic eigensystems so that realistic cloth dynamics can be more efficiently simulated without the need of numerical eigendecomposition and backtracking-based line search filtering. Concurrently, Ando (2024) proposed a cubic barrier energy to strictly enforce penetration-free and strain-limiting constraints, with the aim of avoiding vanishing search steps when the constraint gaps are tiny.
@@ 3. Background and preliminaries
Figure 2. An overview of 1 Newton iteration in our unified GPU IPC framework. Our simulation focuses on cloth, soft bodies with varying stiffness, and stiff affine bodies (Lan et al., 2022a). Cloth and soft bodies are modeled using FEM. Contact between all objects is handled using IPC (Li et al., 2020a). We solve the system with a Newton-PCG solver and use Continuous Collision Detection (CCD) to filter the search direction and prevent penetration, applying backtracking line search to ensure energy decrease. Our contributions, highlighted in light pink blocks, include a novel inexact strain limiting energy with an analytic eigensystem for efficient PSD projection, an optimized linear solver with a connectivity-enhanced MAS preconditioner, and a highly-parallelized global Hessian assembly strategy for affine-deformable coupling.
@@ 3.1. Unified Incremental Potential
where $\mathbf{q}\in\rm I\!R^{12\alpha}$ is the reduced space coordinates of the $\alpha$ affine bodies. For affine body $j$, $\mathbf{q}_{j}=[\mathbf{p}_{j}^{T},\mathbf{A}_{j1}^{T},\mathbf{A}_{j2}^{T},\mathbf{A}_{j3}^{T}]^{T}$, with $\mathbf{p}_{j}\in\rm I\!R^{3}$ the translation vector and $\mathbf{A}_{j}\in\rm I\!R^{3\times 3}$ the affine deformation matrix. Here, $\mathbf{x}_{i}(\mathbf{q}_{j})=\mathbf{A}_{j}\bar{\mathbf{x}}_{i}+\mathbf{p}_{j}$ is the current position of full-space node $i$ on affine body $j$ for measuring contact energies, with $\bar{\mathbf{x}}_{i}$ the rest-state full space coordinates. The Jacobian matrix $\mathbf{J}_{ij}=\frac{\partial\mathbf{x}_{i}}{\partial\mathbf{q}_{j}}\in\rm I\!R^{3\times 12}$ maps information between the reduced and full spaces. For example, the mass matrix $\mathbf{M}_{r}$ of affine body $j$ is calculated as $\mathbf{M}_{r}=\sum_{i}m_{i}\mathbf{J}_{ij}^{T}\mathbf{J}_{ij},$ where $m_{i}$ is the lumped mass corresponding to each node $i$. Similarly, the contact gradient of affine body $j$ is calculated as $\nabla_{\mathbf{q}_{j}}B=\sum_{i}\mathbf{J}_{ij}^{T}\nabla_{\mathbf{x}_{i}}B.$ From this, we can see that computing contact forces and Hessian matrices for the reduced DOFs could require accumulating values from a significant amount of contact pairs, which can be expensive.
@@
To accelerate local Hessian computation on the GPU, we utilize potential energies with analytic eigensystems (subsection 3.2). Specifically, for collision energy $B$ and friction energy $D$, we employ the eigenanalysis from GIPC (Huang et al., 2024). The Stable Neo-Hookean model (Smith et al., 2018) is applied for $\Psi_{vol}$, while the bending energy from (Wu and Kim, 2023) is used for $\Psi_{bend}$. To efficiently assemble $\Psi_{memb}$ and $\Psi_{strain}$, we use Kim (2020)’s membrane energy and propose a cubic inexact strain-limiting energy with analytic eigensystem, detailed in section 5.
@@ 3.2. IP Minimization and Linear Solves
To minimize $E(\mathbf{u})$, where $\mathbf{u}=\{\mathbf{q},\mathbf{x}\}$, line search methods are often used, which compute a descent direction by minimizing a local quadratic proxy of the IP in iteration $i$:
(6) $E_{i}(\mathbf{u})=E(\mathbf{u}_{i})+\left(\mathbf{u}-\mathbf{u}_{i}\right)^{T}\nabla E(\mathbf{u}_{i})+\frac{1}{2}\left(\mathbf{u}-\mathbf{u}_{i}\right)^{T}\mathbf{P}(\mathbf{u}_{i})\left(\mathbf{u}-\mathbf{u}_{i}\right),$
and then search along this direction to find a new iterate. The descent direction is obtained by solving the linear system $\mathbf{P}(\mathbf{u}_{i})\mathbf{d}=-\nabla E(\mathbf{u}_{i})$, where $\mathbf{d}=\mathbf{u}_{i}^{*}-\mathbf{u}_{i}$ and $\mathbf{u}_{i}^{*}=\operatorname*{arg\,min}_{\mathbf{u}}E_{i}(\mathbf{u})$. Here, $\mathbf{P}$ is a symmetric positive definite (SPD) proxy matrix approximating $\nabla^{2}E$ for fast convergence, with projected Newton (Li et al., 2020a) or Gauss-Newton (Huang et al., 2024) approximations shown effective. The new iterate is computed as $\mathbf{u}_{i+1}=\mathbf{u}_{i}+\alpha\mathbf{d}$, where $\alpha$ is the step size calculated via backtracking line search, with filtering to ensure feasibility imposed by contact and strain limits (Li et al., 2021). The minimization terminates when the iterate is sufficiently close to the local minimum, e.g. when $\|\mathbf{d}\|\leq\varepsilon_{d}$, where $\varepsilon_{d}$ is the Newton tolerance.
@@ 4.1. Discussions of MAS
MAS constructs the domain hierarchy by first sorting all mesh vertices using Morton codes, and then grouping these vertices into subdomains of size $N$. Next, connected vertices within each subdomain are merged to create the next-level simulation domain. These merged vertices, called super nodes, are then grouped again to form the domains at subsequent levels until a maximum level is reached, no further merges could be performed, or only a single subdomain is present. During preconditioning, the inverse of each subdomain’s Hessian matrix is multiplied with the mapped input vector, and the resulting vectors are mapped and summed to produce the output (see our supplemental document for more details). With a fixed number of levels, maximizing node connectivity within each subdomain is critical to making dimensionality reduction more effective, which allows information to propagate farther at coarser levels and ensures faster convergence. See our supplemental document for more details.
@@ 4.2. Connectivity-Enhanced MAS Construction
To implement this approach, we first construct a mapping array of size $N\cdot M$, which maps each node to a slot in its corresponding MAS subdomain, as shown in Figure 7. This mapping array is then used to build the MAS preconditioner. To ensure correct data access during construction, we also create a remapping array, which helps GPU threads access the correct node data for each entry in the mapping array. Padded entries, where no nodes are mapped, are marked with the value $-1$ (see Figure 7). The remaining MAS levels are constructed using the original method based on node connectivity. Considering the connectivity introduced by the current contact pairs in each Newton iteration, our approach reduces connectivity issues and enables more effective aggregation, improving the convergence speed of the MAS-preconditioned linear solver by up to 2.4$\times$. Note that the METIS partitioning is not updated during the simulation, as we observed that it does not provide further speedup.